OpenStudy (anonymous):

Find the x and y intercept 3/5x-12

OpenStudy (anonymous):

Can you give the question clearly

OpenStudy (anonymous):

3over5*x-12

OpenStudy (anonymous):

y = mx+b where b is the y intercept. set y equal to zero and solve for x to get the x intercept

OpenStudy (anonymous):

well put x = 0 you get y = -12 put y = 0, you get 3/5x = 12 5x = 1/4 x=1/20

OpenStudy (anonymous):

y intercept = -12 x intercept = 20

OpenStudy (anonymous):

0 = (3/5)x-12 12 = (3/5)x (5/3)*12 = x 60/3 = x x=20

OpenStudy (anonymous):

u messed up somewhere kippendorf

OpenStudy (anonymous):

You re may be deezil_wood but the question posted is not clear.

OpenStudy (anonymous):

yeah it's not so i assumed it was in standard slope-intercept form he retyped it and it's still not clear

OpenStudy (anonymous):

I meant, you may be right lol.

OpenStudy (anonymous):

They should have an edit function to your replies.

OpenStudy (anonymous):

the way he retyped it, now it's 3 / (5x) - 12

OpenStudy (anonymous):

so we're both right lol

OpenStudy (anonymous):

yeah. which would make my answer correct but not possible i don't think